OpenStudy (anonymous):

Make x the subject y = (2x-1)/(x+3)

OpenStudy (anonymous):

\[y=\frac{ 2x-1 }{x+3 }\]

OpenStudy (anonymous):

X= -10,-4,-2,4 Y=3,9,-5,1

OpenStudy (phi):

as a first step, multiply both sides by (x+3) like this: \[ y\cdot (x+3)=\frac{ 2x-1 }{x+3 } \cdot (x+3) \]

OpenStudy (anonymous):

Yes I've done that but I don't know to how continue ...

OpenStudy (phi):

what do you get after simplifying the right side ?

OpenStudy (anonymous):

I solved for x and y, they have more than one solution.

OpenStudy (anonymous):

y(x+3)=2x-1

OpenStudy (anonymous):

make xs in one side and ys the otherside

OpenStudy (anonymous):

We're making x the subject not solving...

OpenStudy (phi):

now distribute the y : multiply each term in the parens by y

OpenStudy (anonymous):

yx + 3y = 2x-1

OpenStudy (phi):

I would add -yx to both sides and add +1 to both sides

OpenStudy (anonymous):

3y+1=2x-yx?

OpenStudy (phi):

almost done... factor out the x from each term on the right side

OpenStudy (phi):

factoring x out is the opposite of "distributing"

OpenStudy (anonymous):

3y+1=x(2-y) x=(3y+1)/(2-y)

OpenStudy (phi):

yes

OpenStudy (anonymous):

Thank you for your help!

OpenStudy (phi):

yw